# TaskCLIP: Extend Large Vision-Language Model for Task Oriented Object Detection

### v1

@@ 3.4 Score function and training
After m transformer aligner layers’ processing, we get the recalibrated image vision embedding ($\textbf{E}^{\textbf{m}}_{\textbf{V}}$) and text embedding ($\textbf{E}^{\textbf{m}}_{\textbf{T}}$). In Fig. 2(a)
5
5
[…]
, we perform matrix to matrix multiplication between $\textbf{E}^{\textbf{m}}_{\textbf{T}}$ and $\textbf{E}^{\textbf{m}}_{\textbf{T}}$ to generate the predicted affinity matrix $\textbf{A}_{p}$. The computation could me summarized as:
@@
The shape of the affinity matrix $\textbf{A}_{\textbf{p}}$ is $N_{bbx}\times N_{word}$. In traditional CLIP, to perform the zero-shot image classification, we will directly apply softmax function over affinity matrix to get the predict label. In task-oriented object detection, since we need to find the most-optimal items, further processing is necessary. Regarding the meaning of $\textbf{A}_{\textbf{p}}$[i][j], it represents the score that bounding box i have visual attribute j. To make the final decision of choosing the most optimal item that could be treated as the ideal task affordance, we pass the affinity matrix into a score function, as is shown in Fig. 2(a)
6
6
[…]
. In Fig. 2(c), we present the score function model architecture. It will first encode each bounding box’s visual feature score vector into a higher dimensional vector. In Fig. 2(c), we represented the encoded new matrix as $\textbf{H}_{\textbf{p}}$ whose dimension is $N_{bbox}\times D^{\prime}$. This encoding process could be represented as:
@@
The we pass $\textbf{H}_{\textbf{p}}$ into a self-attention layer. Here, we incorporate a self-attention layer because we aim to learn the relative importance of different items, enabling us to select the most suitable items for this task. Following the self-attention layers, we employ multiple fully connected layers for regression, ultimately producing the score vector, ($\vec{S}_{p}$). The shape of $\vec{S}_{p}$ is $1\times N_{bbox}$ while each element of $\vec{S}_{p}$ has value ranging from 0 to 1. For inference, we will set a threshold value ($\delta$) and we have:
@@
Here $\textbf{P}[i]$ is the prediction of whether the item of bounding box i is an ideal affordance of this task. During the training, we will have the ground truth score vector $\vec{S}_{g}$. For the $i^{t}h$ element of $\vec{S}_{g}$, if it can solve the task then its value is 1 otherwise its value is 0. The loss L during the back propagation is:
@@
Here $MSELoss$ is the mean squared error loss function [16]. After getting the loss, we perform the back propagation and conduct end to end training of both transformer aligner module and score function as is shown in Fig. 2(a)
8
8
[…]
. Unlike end-end to training style of DETR-based works, our framework will not touch object detection network and vision-language model. Reuse vision-language pretraining knowledge make our framework has higher training efficiency when compared with previous works, as is shown in section 4.
@@ 3.5 Select-by-grouping mechanism
When using existing object detection network, such as Faster R-CNN, we not just get the bounding box (bbox) information but also the bbox COCO class. To assist us find all suitable affordance, here we propose the select-by-grouping mechanism. Suppose in a image, we find total N bbox. After the score function, we find the $i^{th}$ bbox’s regression score is higher than this task’s threshold. Considering the bbox quality, it is possible that we omit some items which are the same COCO class of the $i^{t}h$ bbox. In most case, the same class items show high chance to be ideal task affordance. To compensate this false-negative drawback, our select-by-grouping mechanism will check all other bbox and see if the following scenario satisfied:
$\mathbf{P}[j]=1\quad if\quad\mathbf{class}[j]=\mathbf{class}[i]\quad and\quad\mathbf{conf}[j]>\beta\quad and\quad\mathbf{conf}[i]>\beta$ (7)
@@ 4.1 Datasets and Metric
Table 1: COCO-Tasks dataset task name.
task1: step on task2: sit confortably task3: place flowers task4: get potatoes out of fire
task5: water plant task6: get lemon out of tea task7: dig hole task 8: open bottle of beer
task9: open parcel task10: serve wine task11: pour sugar task12: smear butter
task13: extinguish fire task 14: pound carpet - -
@@ 4.2 Implementation Details
For the object detection network, we experiment with both Faster R-CNN [32] and YOLOv8 [17]. The implementation of Faster R-CNN utilizes the X101-FPN configuration in Detectron2 [38]. As for YOLOv8, we employ the YOLOv8x configuration. Regarding the large vision language model (VLM), we employ OpenCLIP [4, 31]. Specifically, the vision transformer encoder configuration is ViT/H and the text encoder is RoBERTa [25]. We implement transformer aligner module and score function using PyTorch [30]. The model was trained for 20 epochs with an initial learning rate of 1e-6, employing AdamW [26] as the optimizer. We generated visual feature attributes for each task using OpenAI ChatGPT 3.5 [27].
@@
Table 3: Comparison with the state-of-the-art methods.
Task(AP@0.5)(%) mAP@0.5(%)
task1 task2 task3 task4 task5 task6 task7 task8 task9 task10 task11 task12 task13 task14 mean
GGNN [33] 36.6 29.8 40.5 37.6 41.0 17.2 43.6 17.9 21.0 40.6 22.3 28.4 39.1 40.7 32.6 (+0.0)
TOIST [20] 44.0 39.5 46.7 43.1 53.6 23.5 52.8 21.3 23.0 46.3 33.1 41.7 48.1 52.9 41.3 (+8.5)
@@
[…]
TaskCLIP 44.9 44.1 45.4 56.5 50.2 32.3 66.0 19.7 38.0 52.1 37.8 41.6 44.9 64.0 45.5 (+12.9)
TaskCLIP* 45.1 45.1 48.6 59.1 52.9 33.0 67.3 22.1 39.0 55.7 39.9 43.5 46.5 66.3 47.4 (+14.8)
[…]
@@ 4.3 Comparisons with Previous Works
In Table 3, we present a comparison of TaskCLIP’s performance with previous works on the COCO-Task dataset. We utilize Faster R-CNN as the front-end object detection network to identify bounding boxes (bbox) for each item. The transformer aligner module between OpenAI CLIP and the score function comprises 8 layers, with each layer’s self-attention and cross-attention having an attention head size ($N_{Head}$) of 4. We set the threshold for each score function output to 0.15 for all tasks as the baseline TaskCLIP result. Additionally, we provide TaskCLIP’s performance with various optimizations in Table 3. In Section 4.4, we will delve into the effects of the transformer aligner module’s layer size ($N_{Layer}$), attention head size ($N_{Head}$), and threshold value ($\delta$) on the detection accuracy (AP@0.5).
Among the previous works focusing on task-oriented object detection, TOIST and CoTDet are based on DETR, while the work [33] employs a pre-trained object detection network as the front-end. However, in the back-end, work [33] use Gated Graph Neural Network (GGNN) [21] to select objects that could potentially serve as the ideal affordance for specific tasks. TaskCLIP achieves a 4.5% higher mAP@0.5 than the state-of-the-art (SOTA) model, CoTDet [36] (our own implementation version), without employing knowledge distillation. Compared to TOIST (using noun-pronoun distillation), TaskCLIP achieves a 3.3% improvement in mAP@0.5 accuracy. As shown in Table 2, TaskCLIP’s training efficiency surpasses that of TOIST due to TaskCLIP’s utilization of existing pre-trained vision-semantic information rather than recalibrating them from scratch.
@@ 4.4.1 Hyperparameter Tuning
*Here we use blue and red color to mark the top 3 highest mAP@0.5 under threshold 0.2 and threshold 0.4.
In this section, we focus on three hyperparameters: the number of hidden layers of the transformer aligner module ($N_{Layer}$), the number of attention heads in each aligner layer ($N_{Head}$), and the score function threshold ($\delta$). In Table 4, we present the effects of $N_{Layer}$ and $N_{Head}$ on mAP@0.5 across 14 tasks, under two different $\delta$ values, 0.2 and 0.4. We choose Faster R-CNN as the object detection network. As noted by previous studies [33], the occurrence of suitable affordances for each task in the COCO-Tasks dataset is much lower than that of not ideal affordances, with an average ratio of $\sim 1:15$. To address this, we set the threshold below 0.5 for our ablation study of the transformer aligner architecture. Table 4 demonstrates that the top three configurations yielding the best accuracy results are (8,4), (8,16), and (10,16) in the format of ($N_{Layer}$,$N_{Head}$). Among these, we believe the combination ($N_{Layer}$: 8, $N_{Head}$: 4) strikes the best balance between model size and accuracy.
Figure 4: Precision-recall curves of all 14 tasks. For each task, we try three different transformer aligner module configuration based on previous experiment. Inside each task’s PR curve, we also mention the AUC and threshold ($\delta$) as (AUC, $\delta$)
In the investigation of threshold effects, we consider the data imbalance of COCO-Tasks. Here, we employ g-means [15] to determine each task’s optimal threshold. We can generally view the score function component of TaskCLIP as a binary classification problem when using ground truth bounding boxes (gt bbox). Denoting the false positive rate and true positive rate as fpr and tpr respectively, we explore the range of $\delta$ from 0 to 1 with a step size of 0.01. The g-means equation for task $j$ can be expressed as follows:
$g_{j}=\sqrt{tpr\times(1-fpr)}$ (8)
And then we can calculate the best threshold ($\delta$) for the task j:
$\delta=argmax(g_{j})*0.01$ (9)
In Fig. 4, we illustrate each task’s Precision-Recall curve (PR curve) along with the best threshold and Area Under the Curve (AUC) for each task. The average threshold ($\delta$) over all 14 tasks for aligner configuration (8,4) is 0.15, which is also the $\delta$ value used in Table 3.
@@
(b) F-RCNN 44.9 44.1 45.4 56.5 50.2 32.3 66.0 19.7 38.0 52.1 37.8 41.6 44.9 64.0 45.5 (+19.5)
(c) F-RCNN+diff $\delta$ 44.9 43.3 45.1 56.9 50.1 32.7 64.5 20.4 38.2 53.7 37.5 41.9 46.1 63.5 45.6 (+19.6)
(d) F-RCNN+SG 45.5 44.8 46.1 57.8 50.8 32.9 66.5 20.3 38.3 52.8 38.5 41.6 45.6 65.5 46.2 (+20.2)
(e) F-RCNN+both 45.1 45.1 48.6 59.1 52.9 33.0 67.3 22.1 39.0 55.7 39.9 43.5 46.5 66.3 47.4 (+21.4)
(f) YOLOv8 48.7 46.6 45.6 57.1 50.1 30.9 62.5 22.0 38.0 53.9 36.9 40.0 42.6 63.4 45.5 (+19.5)
(g) YOLOv8+diff $\delta$ 48.7 45.4 45.6 57.1 50.1 31.1 61.0 23.1 38.2 55.6 36.3 40.5 43.1 63.3 45.7 (+19.7)
(h) YOLOv8+SG 48.6 46.6 45.9 57.1 50.4 31.1 62.5 22.2 38.3 54.3 37.1 40.0 42.6 64.4 45.8 (+19.8)
(i) YOLOv8+both 49.0 46.2 46.0 57.5 50.4 31.3 61.5 23.2 38.5 55.6 36.6 40.5 43.3 63.9 46.0 (+20.0)
(j) ground truth bounding box 61.0 72.8 56.1 76.4 56.4 47.2 81.3 28.5 49.7 67.5 47.7 55.6 51.9 81.3 59.5
@@ 4.4.2 TaskCLIP Component Analysis
Table 5 illustrates the contribution of each component of TaskCLIP to the final accuracy. These components include object detection network type, transformer aligner module, and select-by-grouping mechanism. We compare the performance with the baseline, which directly selects the objects based on the affinity matrix without any re-calibration and alignment.
For assessing the effect of the transformer aligner, we present the accuracy of the combination of OpenCLIP and the score function ($f_{score}$). The original OpenAI CLIP and OpenCLIP models were primarily designed to match images with nouns, making it challenging for them to accurately match each bounding box item image with its corresponding visual feature attributes. As depicted in Table 5 (c) and (g), the mAP@0.5 improves by approximately 20%.
For Table 5 (b-e) and Table 5 (f-i), we depict the effects of various optimizations on TaskCLIP’s final performance. Specifically, Table 5 (b-e) utilizes Faster R-CNN, while Table 5 (f-i) employs YOLOv8. In Table 5 (j), we present the AP@0.5 results obtained when using the ground truth bounding boxes. We conduct two pairs of comparisons, namely (b) and (c), (f) and (g), to illustrate the effect of the different threshold strategy (diff $\delta$). In (b) and (f), we utilize the average threshold of 0.15, as reported in Fig. 4, while in (c) and (g), we apply different thresholds based on g-means for each task. The accuracy improvement achieved with different $\delta$ values is 0.1% and 0.2% when using Faster R-CNN and YOLOv8 as the object detection network, respectively.
In Table 5.(d) and Table 5.(h), we present the accuracy improvement when using select-by-grouping mechanism ($SG$). Compared to normal setting, $SG$ mechanism provide 0.7% and 0.3% accuracy improvement when using Faster R-CNN and YOLOv8 as the detection network.
In Table 5 (e) and Table 5 (i), we present the final AP@0.5 results obtained when deploying both optimization techniques. With the implementation of both optimizations, TaskCLIP achieves a 21.4% and 20.0% improvement in mAP@0.5 compared to the baseline (Table 5 (a)). TaskCLIP effectively recalibrates the item embeddings (vision space) with task-related attribute embeddings (text embedding). However, when compared to using ground truth bounding boxes (Table 5 (j)), TaskCLIP experiences approximately a 12% decrease in mAP@0.5. We attribute this drop in accuracy primarily to limitations inherent in the detection network. As we will elaborate in section 4.5, the bounding boxes generated by Faster R-CNN and YOLOv8 exhibit deviations when compared to the ground truth bounding boxes.

### v2

@@ 3.5 Score function and training
After m transformer aligner layers’ processing, we get the recalibrated image vision embedding ($\textbf{E}^{\textbf{m}}_{\textbf{V}}$) and text embedding ($\textbf{E}^{\textbf{m}}_{\textbf{T}}$). In Fig. 3(a), we perform matrix to matrix multiplication between $\textbf{E}^{\textbf{m}}_{\textbf{T}}$ and $\textbf{E}^{\textbf{m}}_{\textbf{T}}$ to generate the predicted affinity matrix $\textbf{A}_{p}$. The computation could be summarized as:
@@
The shape of the affinity matrix $\textbf{A}_{\textbf{p}}$ is $N_{bbx}\times N_{word}$. In traditional CLIP, to perform the zero-shot image classification, we will directly apply softmax function over affinity matrix to get the predict label. In task-oriented object detection, since we need to find the most-optimal items, further processing is necessary. Regarding the meaning of $\textbf{A}_{\textbf{p}}$[i][j], it represents the score that bounding box i have visual attribute j. To make the final decision of choosing the most optimal item that could be treated as the ideal task affordance, we pass the affinity matrix into a score function. In Fig. 3(c), we present the score function model architecture. It will first encode each bounding box’s visual feature score vector into a higher dimensional vector. In Fig. 3(c), we represented the encoded new matrix as $\textbf{H}_{\textbf{p}}$ whose dimension is $N_{bbox}\times D^{\prime}$. This encoding process is:
@@
Then we pass $\textbf{H}_{\textbf{p}}$ into a self-attention layer. Here, we incorporate a self-attention layer because we aim to learn the relative importance of different items, enabling us to select the most suitable items for this task. Following the self-attention layers, we employ multiple fully connected layers for regression, ultimately producing the score vector, ($\vec{S}_{p}$). The shape of $\vec{S}_{p}$ is $1\times N_{bbox}$ while each element of $\vec{S}_{p}$ has value ranging from 0 to 1. For inference, we will set a threshold value ($\delta$) and we have:
@@
Here $\textbf{P}[i]$ is the prediction of whether the item of bounding box i is an ideal affordance of this task. During training, we will have the ground truth score vector $\vec{S}_{g}$. For the $i^{t}h$ element of $\vec{S}_{g}$, if it can solve the task then its value is 1 otherwise its value is 0. The loss L during the back propagation is:
@@
Here $MSELoss$ is the mean squared error loss function [19]. After getting the loss, we perform the back propagation and conduct end to end training of CLIP adapter, global attention layer, transformer aligner module, and score function as is shown in Fig. 3(a). Unlike end-end to training style of DETR-based works, our framework will not touch object detection network and vision-language model. Reuse vision-language pretraining knowledge make our framework has higher training efficiency when compared with previous works, as is shown in section 4.
Table 2: Original COCO-Tasks dataset task name and corresponding synonym task name from ChatGPT. The synonym task name is in the parentheses.
task1: step on (trend on) task2: sit comfortably (lounge easily) task3: place flowers (arrange blooms)
task4: get potatoes out of fire (retrieve tubers from heat) task5: water plant (irrigate vegetation) task6: get lemon out of tea (remove citrus from brew)
task7: dig hole (excavate pit) task 8: open bottle of beer (uncap ale) task9: open parcel (unwrap package)
task10: serve wine (serve a cocktail) task11: pour sugar (dispense sweetener) task12: smear butter (spread cream)
task13: extinguish fire (douse flames) task 14: pound carpet (beat rug)
@@ 3.6 Select-by-grouping mechanism
When using existing object detection network, such as Faster R-CNN, we not just get the bounding box (bbox) information but also the bbox COCO class. To assist us find all suitable affordance, here we propose the select-by-grouping mechanism. Suppose in an image, we find a total of N bbox. After the score function, we find that the $i^{th}$ bbox’s regression score is higher than this task’s threshold. Considering the bbox quality, it is possible that we omit some items which are the same COCO class of the $i^{t}h$ bbox. In most cases, the same class items show a high chance to be ideal task affordance. To compensate this false-negative drawback, our select-by-grouping mechanism will check all other bbox and see if the following scenario satisfied:
$\begin{split}\mathbf{P}[j]=1\quad&\text{if}\quad\mathbf{class}[j]=\mathbf{class}[i]\\ &\text{and}\quad\mathbf{conf}[j]>\beta\\ &\text{and}\quad\mathbf{conf}[i]>\beta\end{split}$ (9)
@@ 4.1 Datasets and Metric
Task-agnostic experiment To evaluate the model’s ability to generalize, we not only used the 14 original tasks from the COCO-Tasks dataset but also performed a task-agnostic experiment. Initially, referencing Tab. 1.(b), for each original task in the COCO-Tasks dataset, we requested ChatGPT to generate synonymous task words, as displayed in Tab. 2. For instance, the synonym task for Serve wine? is Serve a cocktail?. These two tasks share common affordances (such as wine glasses), so an effective generalization model trained on the original COCO-Tasks dataset should be capable of addressing the new task.
@@ 4.2 Implementation Details
For the object detection network, we experiment with both Faster R-CNN [40] and YOLOv8 [22]. The implementation of Faster R-CNN utilizes the X101-FPN configuration in Detectron2 [47]. As for YOLOv8, we employ the YOLOv8x configuration. Regarding the large vision language model (VLM), we employ OpenCLIP [4, 38]. Specifically, the vision transformer encoder configuration is ViT/H and the text encoder is RoBERTa [31]. We implement transformer aligner module and score function using PyTorch [37]. The model was trained for 20 epochs with an initial learning rate of 1e-6, employing AdamW [32] as the optimizer. We generated visual feature attributes for each task using OpenAI ChatGPT [33].
@@
Table 4: Comparison with previous methods. Values in parentheses indicate results from task-agnostic experiments.
@@
task1 36.6 (-) 44.0 (-) 45.8 (-) 58.9 (56.2) 48.4 (47.6) 52.1 (50.9)
task2 29.8 (-) 39.5 (-) 40.0 (-) 55.0 (53.1) 44.6 (44.1) 47.4 (46.6)
task3 40.5 (-) 46.7 (-) 49.4 (-) 51.2 (48.2) 52.4 (52.4) 54.1 (54.1)
task4 37.6 (-) 43.1 (-) 49.6 (-) 68.5 (57.3) 60.6 (60.6) 62.5 (62.5)
task5 41.0 (-) 53.6 (-) 53.4 (-) 60.5 (54.9) 54.4 (54.4) 56.9 (56.9)
task6 17.2 (-) 23.5 (-) 26.9 (-) 47.7 (35.0) 31.5 (31.5) 34.0 (34.0)
task7 43.6 (-) 52.8 (-) 58.3 (-) 76.9 (50.5) 67.7 (67.7) 69.9 (69.9)
task8 17.9 (-) 21.3 (-) 22.6 (-) 40.7 (3.6) 18.6 (18.6) 19.2 (19.2)
task9 21.0 (-) 23.0 (-) 32.5 (-) 47.4 (20.1) 39.4 (39.4) 41.4 (41.4)
task10 40.6 (-) 46.3 (-) 50.0 (-) 66.5 (58.5) 55.7 (55.7) 58.1 (58.1)
task11 22.3 (-) 33.1 (-) 35.5 (-) 41.9 (30.6) 38.1 (38.1) 39.8 (39.8)
task12 28.4 (-) 41.7 (-) 43.7 (-) 48.3 (47.3) 47.6 (47.6) 49.8 (49.8)
task13 39.1 (-) 48.1 (-) 52.8 (-) 61.7 (9.4) 46.3 (46.3) 48.7 (48.7)
task14 40.7 (-) 52.9 (-) 56.2 (-) 71.4 (58.4) 67.4 (67.4) 69.8 (69.8)
mean (mAP@0.5(%)) 32.6 (-) 41.3 (-) 44.1 (-) 56.9 (40.9) 48.1 (47.9) 50.3 (50.1)
@@ 4.3 Comparisons with Previous Works
In Table 4, we present a comparison of TaskCLIP’s performance with previous works on the COCO-Task dataset. Here we present results for both original COCO-Task dataset tasks’ results and synonym tasks’ result (in the parentheses). We utilize Faster R-CNN as the front-end object detection network to identify bounding boxes (bbox) for each item. The transformer aligner module between OpenAI CLIP and the score function comprises 8 layers, with each layer’s self-attention and cross-attention having an attention head size ($N_{Head}$) of 4. We set the threshold for each score function output to 0.15 for all tasks as the baseline TaskCLIP result. Additionally, we provide TaskCLIP’s performance with various optimizations in Table 4. In Section 4.4, we will delve into the effects of the transformer aligner module’s layer size ($N_{Layer}$), attention head size ($N_{Head}$), and threshold value ($\delta$) on the detection accuracy (AP@0.5).
Among the previous works focusing on task-oriented object detection, TOIST and CoTDet are based on DETR, while the work [41] employs a pre-trained object detection network as the front-end. However, in the back-end, work [41] use Gated Graph Neural Network (GGNN) [27] to select objects that could potentially serve as the ideal affordance for specific tasks. Given that both TOIST and GGNN require human involvement to prepare words for specific tasks, they are task-dependent in an unnatural manner, making it challenging to apply them to other tasks. Therefore, for a task-agnostic evaluation, we only consider CoTDet and TaskCLIP. To assess CoTDet’s performance on the synonym task, we adhere to its ChatGPT prompt and utilize RoBERTa to produce new item and visual feature text embeddings. For TaskCLIP, we reference Tab. 1.(c), requesting ChatGPT to choose visual description words from the visual description word banks maintained during the training phase. These newly selected description words are then directly used to enable TaskCLIP to perform object detection.
As indicated by Tab. 4, TaskCLIP surpasses both GGNN and TOIST on the original COCO-Tasks dataset. Although TaskCLIP’s accuracy is lower than CoTDet’s on the original COCO-Tasks dataset, it outperforms CoTDet by over 9% on the new synonyms dataset. This is because CoTDet’s training involves direct training of descriptive words and affordance text embedding vectors. Consequently, the accuracy of CoTDet significantly declines when new tasks arise and ChatGPT changes its responses. Conversely, TaskCLIP concentrates on learning the visual features of tasks, which results in better generalizability.
As shown in Table 3, TaskCLIP’s training efficiency surpasses that of TOIST and CoTDet due to TaskCLIP’s utilization of existing pre-trained vision-semantic information rather than recalibrating them from scratch.
@@ 4.4.1 Hyperparameter Tuning
In this section, we focus on three hyperparameters: the number of hidden layers of the transformer aligner module ($N_{Layer}$), the number of attention heads in each aligner layer ($N_{Head}$), and the score function threshold ($\delta$). In Table 5, we present the effects of $N_{Layer}$ and $N_{Head}$ on mAP@0.5 across 14 tasks, under two different $\delta$ values, 0.2 and 0.4. We choose Faster R-CNN as the object detection network. As noted by previous studies [41], the occurrence of suitable affordances for each task in the COCO-Tasks dataset is much lower than that of not ideal affordances, with an average ratio of $\sim 1:15$. To address this, we set the threshold below 0.5 for our ablation study of the transformer aligner architecture. Table 5 demonstrates that the top three configurations yielding the best accuracy results are (8,4), (8,16), and (10,16) in the format of ($N_{Layer}$,$N_{Head}$). Among these, we believe the combination ($N_{Layer}$: 8, $N_{Head}$: 4) strikes the best balance between model size and accuracy. In the investigation of threshold effects, we consider the data imbalance of COCO-Tasks. Here, we employ g-means [18] to determine each task’s optimal threshold. The average threshold ($\delta$) over all 14 tasks for aligner configuration (8,4) is 0.15, which is also the $\delta$ value used in Table 4.
@@
(b) F-RCNN 48.4 44.6 52.4 60.6 54.4 31.5 67.7 18.6 39.4 55.7 38.1 47.6 46.3 67.4 48.1 (+22.1)
(c) F-RCNN+diff $\delta$ 48.4 44.8 52.4 60.6 54.5 31.5 67.6 18.8 39.4 55.9 38.1 47.6 46.3 67.4 48.6 (+22.6)
(d) F-RCNN+SG 52.0 47.4 54.1 62.5 56.7 34.0 69.9 19.2 41.5 58.1 39.8 49.8 48.7 69.8 50.1 (+24.1)
(e) F-RCNN+both 52.1 47.4 54.1 62.5 56.9 34.0 69.9 19.2 41.4 58.1 39.8 49.8 48.7 69.8 50.3 (+24.3)
(f) YOLOv8 48.7 46.6 45.6 57.1 50.1 30.9 62.5 22.0 38.0 53.9 36.9 40.0 42.6 65.4 47.5 (+21.5)
(g) YOLOv8+diff $\delta$ 48.7 45.4 45.6 57.1 50.1 31.1 61.0 23.1 38.2 55.6 36.3 40.5 43.1 65.3 47.7 (+21.7)
(h) YOLOv8+SG 48.6 46.6 45.9 57.1 50.4 31.1 62.5 22.2 38.3 54.3 37.1 40.0 42.6 64.4 47.8 (+21.8)
(i) YOLOv8+both 49.0 46.2 46.0 57.5 50.4 31.3 61.5 23.2 38.5 55.6 36.6 40.5 43.3 66.9 49.0 +(23.0)
(j) ground truth bounding box 61.2 73.0 56.3 76.6 56.6 47.4 81.5 28.7 49.9 67.7 47.9 56.0 52.1 81.5 62.5
@@ 4.4.2 TaskCLIP Component Analysis
Table 6 illustrates the contribution of each component of TaskCLIP to the final accuracy, including the object detection network, transformer aligner module, and select-by-grouping mechanism. We compare the performance with the baseline, which directly selects the objects based on the affinity matrix without any re-calibration and alignment. For assessing the effect of the transformer aligner, we present the accuracy of the combination of OpenCLIP and the score function ($f_{score}$). The original OpenAI CLIP and OpenCLIP models were primarily designed to match images with nouns, making it challenging for them to accurately match each bounding box item image with its corresponding visual feature attributes. For Table 6 (b-e) and Table 6 (f-i), we depict the effects of various optimizations on TaskCLIP’s final performance. TaskCLIP effectively recalibrates the item embeddings (vision space) with task-related attribute embeddings (text embedding). We report the accuracy obtained with ground truth bounding boxes in Table 6 (j).